# DECO: Life-Cycle Management of Enterprise-Grade Chatbots

### v1

@@ 4.1. IcM Processor
Raw information about historical incidents is often scattered across multiple tables in the incident management database. The goal of this module is to extract all relevant details, condense the useful information into a concise and structured format, and create searchable files that can serve as supplementary documentation for the LLM’s context in the backend.
Figure 4 illustrates the pipeline for processing incident management data. The Kusto Connector uses secured authentication through managed identity (Microsoft, 2024m) to access the IcM database and extracts free-form records based on user-specified team name(s) and data range from several data tables. The IcM Processor module uses GPT-4o models to generate structured JSON format summaries from free-form text data based on predefined summarization prompts. Typically, we observe each chatbot creates indexes of around 3000 to 30,000 incidents.
The IcM Processor also functions as an offline platform to “append” additional retrieval fields to documents for the backend module, such as a “helpfulness” score reflecting the thoroughness of the documented investigation process, or “property” fields for specific lookups based on database IDs (see Section 6.2). Ticket summaries and embeddings are stored in a JSON file within Azure Blob Storage, and the IcM index manager initiates the corresponding indexer (Microsoft, 2024i) and index (Microsoft, 2024e), which reference the blob path. A REST API powered by Azure AI Search is deployed to facilitate efficient searches among these files for context-retrieval skills in the backend, using various search algorithms (see Section 5.2).
@@ 4.2. TSG Processor
For TSGs (or any text files), we utilize the original content directly to generate embeddings, splitting them into smaller chunks as needed. To process the documentation, users specify the relevant Git repositories in configuration files, along with corresponding descriptions. These descriptions are included with the retrieved content to provide additional context to the LLM in the backend. Each entry in the configuration involves cloning the repository (using the same authentication mechanism as the IcM processor), selecting specific folders for processing, chunking the files, and generating embeddings for multiple fields, such as document content, titles, and URLs, to enable multi-field hybrid search (see Section 6). Images are extracted in base64 format. The processed files are stored in Blob Storage, and indexed for efficient retrieval by Azure AI Search (similar to IcM as in Figure 4). Additionally, offline preprocessing can generate advanced searchable fields to minimize retrieval latency for the backend. This step supports advanced search algorithms, such as those outlined in (Anonymous, 2024a, b), and Reverse-HyDE question embeddings as in (Gao et al., 2022).
4.3. Code
To create a search index for code, we employ an internal tool at Microsoft, built on the foundation of Tree Sitter (Brunsfeld et al., 2022). This tool constructs a comprehensive Abstract Syntax Tree (AST) to extract key components from the code. These components include classes, functions, statements, variables, and their relationships, such as used by, uses, and inherits from.
The process of constructing the search index involves the following steps: (1) Code Chunking: We generate code chunks based on the structure of the Abstract Syntax Tree (AST). For languages not supported by Tree-sitter, such as Scope (Zhu et al., 2021), we concatenate the code chunk with five preceding and five succeeding chunks. This approach leverages a Large Language Model (LLM) to extract the complete code snippet, ensuring that a single function or class is not split across multiple chunks. The same LLM is also used to identify referenced classes or functions. (2) Component Relationships: For each component, retrieve one-hop related components and attach their corresponding code chunks. (3) Description Augmentation: Use a large language model (LLM) to augment the description of each code chunk. This includes generating a title, a description of the code, references used within the code chunk, and brief descriptions of referenced functions or classes to serve as potential future search fields. (4) Embedding Generation: Produce embeddings for the concatenated code chunks and all associated search fields, establishing a search index given the search fields.
This process creates a robust and flexible search index, enhancing the utility of code components and their contextual relationships for various search scenarios (see Section 6.2).
Figure 5. Overview of PromptFlow modules
@@ 5. Copilot Backend
The backend REST API is implemented using PromptFlow (Microsoft, 2024n) as a stateless API, requiring all necessary information, such as the full chat history as input. This design abstracts away much of the engineering overhead related to authentication and state management, which is handled by our bot frontend (discussed in Section 7). The PromptFlow REST API is composed of various interconnected modules, where each module can be an LLM prompt or a Python script. Figure 5 illustrates the different modules implemented:
  • Skills and Planners: These utilize the Semantic Kernel (Microsoft, 2024j) to generate execution plans based on available skills. Examples include the pre-chat and post-chat skills.
  • Chat Module: This module is responsible for making an LLM call with all the gathered contexts.
  • Auxiliary Modules: These modules support the flow by initializing parameters or managing memory. Examples include the Memory Extractor module and the Prompt Constructor module, which help in truncating or formatting memory as needed.
The search indexes created in Section 4 are utilized by corresponding pre-chat skills to support context retrieval. Most team-specific internal toolkits are implemented as “customized skills” that can be invoked by the pre-chat planner.
5.1. Input and Output
The flow’s input consists of several components: (1) chat history; (2) skill data, typically generated from a dedicated popup UI in the frontend to enforce the invocation of certain skills (see Section 7); (3) the user question; and (4) the user ID, obtained from the frontend (see Figure 5). The flow’s output includes: (1) an answer to be presented to the user; (2) each skill’s output in dictionary format; (3) outputs from the planners (see Section 5.2), indicating which skills to invoke and their corresponding arguments; (4) images generated by the chat for user presentation; and (5) outputs from context-retrieval skills, which are further used for feedback learning and evaluation (see Section 8).
5.2. Skills and Planners
The flow design is centered on supporting the execution of various skills. We initially used the Semantic Kernel (Microsoft, 2024j) sequential planner (without Azure OpenAI function calling (Microsoft, 2024g)) for skill orchestration.
Skill-Chat-Skill Sequence. To minimize the number of calls to large language models (LLMs) and reduce latency, we aim to limit the use of LLMs for most skills. Instead, we combine outputs from various skills and use them jointly as context within the final LLM call. This streamlined approach structures the workflow into a manageable “skill-chat-skill” sequence, reducing the latency of conversations by making only one LLM call. Skills are categorized into two types: those primarily supporting content retrieval (pre-chat) and those that trigger actions post-LLM.
Hierarchical Planners. We observed the following:
Observation: The Semantic Kernel planner struggles as the number of skills and their dependencies increase.
To address system complexity, we implemented hierarchical planners, each managing a subset of skills and incorporating an option to enable Azure OpenAI function calling (Microsoft, 2024g). This approach reduces the complexity faced by each individual planner.
The first group of (pre-chat) skills, referred to as default skills, comprises content retrieval skills that are essential for addressing most inquiries. This group includes skills such as get_code, get_icm, and get_tsg, all backed by Azure AI search indexes.
The second group consists of (pre-chat) customized skills tailored for specific types of queries using team-specific internal tools. An example is a skill that directs users to monitoring dashboards relevant to their particular team. These skills, which are invoked less frequently and are specific to certain teams, can also bypass the LLM chat module, thereby reducing latency (see Figure 5).
The third group of (post-chat) skills is executed after the chat module, either because (1) they depend on outputs from the chat module’s answer (e.g., certain query generation skills with the capability to extract queries from the provided answers are activated only if the chat output includes such queries); or (2) they initiate post-chat actions, such as recommending follow-up questions.
Each skill group is managed by a dedicated planner and an executor (see Figure 5). This organizational structure significantly reduces the planner’s task complexity compared to systems that require detailed logic for comprehensive execution plans (e.g., (Patil et al., 2023; Roy et al., 2024; Song et al., 2023; Yao et al., 2022)).
Skill Dependency. DECO incorporates inter-skill dependencies, allowing multiple skills to use outputs from other skills for contextual grounding. These upstream dependencies are explicitly declared in the skills’ properties. During the execution, skills that do not have dependencies are run concurrently to minimize latency. The execution order, rather than being generated by the planner, is deterministically determined based on the properties of the skills.
5.3. Chat Module
The flow consists of a single chat module, which assembles its prompt template from the following elements: (1) User-specified Prompt: Derived from the prompt_store module using a fine-tuned template, enhanced with team-specific descriptions (see Figure 2) and additional skill directives. (2) Retrieved Documents/Data: Includes skill outputs such as TSG, IcM, code chunks or telemetry data. (3) Current IcM: Provided by the icm_extractor skill when an incident ID is available, providing relevant ticket information for context during the session. (4) Chat History: Integrates previous interactions to ensure continuity. (6) Rephrased User Question: The planner reformulates the user’s question, recorded as user_intent, to enhance search precision. This is particularly useful for contextualizing the query or clarifying the intent. (6) User’s Raw Question: The direct query from the user.
5.4. Auxiliary Modules
The Init module sets environment variables from the user config file, such as the LLM API endpoint, those variables can be then used throughout the flow. The Memory Extractor module truncates the input chat history based on configured parameters of max_chat_history. In the future, this module can help leverage more advanced memory management techniques (such as MemGPT (Packer et al., 2023)) to retrieve/delete relevant information from the full chat history. The Prompt Constructor puts together the user input (such as product description) and merges it with the DECO’s prompt templates and sends the full set of messages to the LLM module, i.e., the Chat module. The skill Validator executes the validation function defined in the class of the invoked skills. For instance, the skill of KQL Generator will validate the generated Kusto queries using LLM to ensure the query is grounded based on retrieved documents.
6. Algorithms
Based on the end-to-end flow design, we implemented several enhancements to further improve chat quality. After analyzing user feedback and bot responses, we observed the following:
Observation: The primary issue with the bot is not hallucination but retrieving incorrect documents or invoking the wrong skills.
This issue likely stems from prompts instructing the bot to rely strictly on retrieved documents rather than generating answers from its own knowledge. While this approach prevents hallucination, it can lead to unhelpful responses like “Sorry, I cannot find relevant information” when incorrect documents or skills are involved. To address this, we implemented key improvements:
  • Planner Improvements: Improved the identification of accurate information sources and tools to ensure the planner selects the appropriate skills for user queries.
  • Retrieval Algorithm Improvements: As opposed to the naive RAG mechanism using a single vector search, we proposed optimized retrieval strategies and reranking algorithms for skills such as get_icm, get_tsg, and get_code to enhance retrieval precision.
These enhancements are supported by an accuracy evaluation system, which forms the basis for the continuous evolution of these algorithms, as discussed in Section 8.
6.1. Planner Improvement
We implemented decorators for each skill to specify its functionality and input arguments, along with examples of positive and negative applicability (Microsoft, 2024j). The planner leverages these annotations to craft execution plans tailored to the available skills. Further enhancements to this framework include:
Few-Shot Examples. Incorporating few-shot examples into the Semantic Kernel enhances the planner’s skill selection accuracy (Microsoft, 2024j) (see Section 8). These examples are included in the prompt argument to refine the execution plan.
Integration with Internal Tools. DECO expands its capabilities by integrating internal tools as customized skills (see Figure 5), including the kql_generator for generating Kusto queries. Kusto is widely used within Microsoft as the primary backend database for storing logs of Azure products and facilitates efficient data management and retrieval. When this skill is executed successfully, it informs the flow execution to bypass the chat module, thus omitting stages like prompt construction and directly presenting the skill’s output. This process significantly accelerates response times for specific queries.
Figure 6. IcM retrieval process
6.2. IcM Retrieval Improvement
@@
The get_icm skill, essential for generating curated incident reports, is frequently used to respond to user inquiries about how-to guides and to search for similar incidents, thereby offering additional mitigation strategies. We’ve noticed that user queries typically highlight specific incident details, such as failure mechanisms or server names, which align with the mitigation steps or extractable properties of incidents (discussed in Section 4). This insight has led us to enhance our retrieval algorithm by implementing detailed filtering of incident candidates using the Azure AI Search API.
Starting with the user’s query, we initiate an “NL2SearchQuery” task using predefined templates (see Figure 6). Depending on the specific input, Azure AI Search enables us to execute tailored search strategies. Each strategy is converted into a search query, executed to gather incidents, and the results are reranked based on a composite score to identify the top $K$ documents for the user.
NL2SearchQuery. For this task, we leverage the default planner’s ability to populate input parameters with values extracted from the chat context (see Figure 5). The planner extracts several key arguments for this skill, including:
@@
  • Search Method: Options include hybrid search, semantic search, simple search, or vector search (Microsoft, 2024h).
@@
Reranking. The initial search query typically returns a large set of incidents, for example, 20. During the reranking step, we calculate scores for each incident to determine the top $K$ results, such as the top 4, for the final retrieval output.
The reranking score for a document $d$ is calculated as follows:
@@
where IS represents the information score, which evaluates the quality of the IcM summary based on data quality metrics like token length, pre-computed during the preprocessing pipeline (see Section 4). TS, the time score, assesses the relevance of incidents based on their age, with older incidents presumed less relevant and assigned lower values. SS, the source score, verifies if the retrieved data aligns with the current context, assigning a value of 1 for matches (e.g., matching team or monitor ID in property fields) and 0 otherwise. These scores are normalized and combined to rerank the retrieved incidents, focusing on factors like information richness, timeliness, and source relevance. This process is detailed in Algorithm 1, and Section 8.2 compares the accuracy of our IcM retrieval algorithm to the existing default similarity search in the IcM Portal.
1:  Input: User question $q$, IcM retrieval database $D$, number of top results $K$, number of incidents to retrieve per query $N$, reranking score function $P(d)$
@@
7:     Compute the reranking score $P(d)$ for all $d\in D_{i}$
@@
10:  Rank all documents in $D^{*}$ based on the reranking score
@@
A similar retrieval mechanism is also employed in the get_code skill for code search. This skill retrieves relevant code snippets in response to user queries by identifying optimal search fields, using output arguments from the Semantic Kernel. The parameters for constructing search queries include:

### v2

@@ 4.2. IcM Processor
The IcM Processor extracts valuable insights from previously mitigated incidents, where engineers may have logged useful queries, mitigation steps, and commands. To better leverage this scattered data, we have decided to preprocess the information into a structured format, formulating additional documentation for the copilot. The output of this pipeline supports the get_icm skill.
Built on the common infrastructure described in Section 4.1, The IcM Processor consists of the following customized steps:
  • Data Injection: The Kusto Connector utilizes secure authentication via managed identity (Microsoft, 2024j) to access the IcM database. It extracts free-form records based on user-specified team name(s) and date range from multiple data tables. The set of IcMs to be included can be customized by different product teams.
  • Data Transformation: The transformation module employs GPT-4o models to generate structured summaries from free-form text data. This process leverages predefined summarization prompts and curated output templates to extract key fields such as summary, title, mitigation steps, and properties to enable specific lookups (see Section 6.2). Previous work has demonstrated that LLMs have effectively tackled this task (Pu et al., 2023).
  • Metadata Enrichment: Additional metadata is computed, such as a helpfulness score, reflecting the thoroughness of the generated documentation.
Figure 3 illustrates the input and output of IcM Processor. Typically, each copilot creates indexes for approximately 3,000 to 30,000 incidents.
Figure 3. Example of data preprocessing pipeline for IcM Processor
@@ 4.3. TSG Processor
The TSG Processor accesses Git repositories, splits long documents into smaller chunks, and generates embeddings for multiple fields, including document content, titles, and URLs. This enables multi-field hybrid search, improving accuracy and ranking. Additional metadata, such as document descriptions, referenced issue IDs, and base64-encoded images are also extracted. The pipeline supports advanced search algorithms, including Reverse-HyDE question embeddings (Gao et al., 2022) and techniques from (Anonymous, 2024a). We conducted a thorough evaluation of documentation retrieval algorithms and hyper-parameters for DECO in (Anonymous, 2024b).
Typically, each copilot indexes approximately 5,000 to 50,000 troubleshooting guide chunks for optimized retrieval.
4.4. Code Processor
Modern software codebases are inherently multi-language, often integrating diverse programming paradigms and a variety of file formats. Beyond source code, typical codebases include configuration files such as YAML, JSON, and XML, which play a critical role in defining system behavior but remain outside the main compilation process. This heterogeneity poses significant challenges for compiler-based indexing tools attempting to comprehensively analyze and query codebases.
Abstract Syntax Tree (AST)-based indexing methods, such as those leveraging Tree-Sitter (Brunsfeld et al., 2022), primarily focus on the syntax of individual programming languages while neglecting other essential components like configuration files, external dependencies, and runtime interactions. As a result, these methods often miss crucial relationships that span across language boundaries. Additionally, because compilation or interpretation usually occurs independently for each language (e.g., Python and C# components are processed separately), traditional indexing techniques struggle to capture the holistic behavior of multi-language systems. Additionally, conventional document chunking and vector-based retrieval approaches frequently struggle to preserve the intricate dependencies and relationships within a codebase. Even worse, these methods risk fragmenting single code segment into multiple chunks, disrupting the continuity of processed documentation and impairing the model’s ability to accurately answer complex user queries.
To address these limitations, DECO introduces the Language-Agnostic RAG (LA-RAG), which leverages LLMs’ multi-language capabilities without relying on compilation. Unlike in-memory Copilot solutions that process only open files within an Integrated Development Environment (IDE), our offline indexing mechanism enables comprehensive search across large-scale codebases containing thousands of files and millions of lines of code spanning multiple languages by structuring the codebase as a loosely coupled reference graph. For the online retrieval, our system is able to accurately determine code construct usage, ensuring context-aware retrieval beyond immediate source files.
The preprocessing pipeline is developed using the same infrastructure as discussed in Section 4. Similarly as the TSG Processor, the Code Processor access the code based from Git repositories, and tentatively splits long code files into fix-size smaller chunks. After generating code chunks, we apply the following metadata enrichment and embedding steps:
  • LLM Rechunking: To prevent functions or classes from being split across chunks, we concatenate each chunk with its five preceding and succeeding chunks, then use an LLM to extract complete code segments.
  • Component Relationships: We use LLM to identify one-hop related components and attach their corresponding code chunks, improving contextual understanding.
  • Description Augmentation: The same LLM call generates a title, a concise description, references used within the code chunk, and summaries of referenced functions or classes to enhance searchability.
  • Embedding Generation: We generate embeddings for code chunks and associated metadata, creating a structured search index for efficient retrieval across multiple fields that is used in the online phase (see Section 6).
Typically, each copilot indexes approximately 300 to 20,000 code chunks for optimized search and retrieval.
@@ 5. Copilot Backend
The backend REST API is implemented as a stateless API, requiring all necessary information, such as the full chat history as input. This design abstracts away much of the engineering overhead related to authentication and state/memory management, which is handled by our bot frontend (discussed in Section 7). The backend orchestration is created to maximize re-usability across different copilots and allow easy customization of incorporate different skills/configuration settings from different teams.
5.1. Design Principles
The DECO’s backend orchestration is guided by the following core principles:
Multi-turn interactions: The future of generative AI platforms lies in multi-agent, multi-turn copilots (Wu et al., 2023). This pattern is emerging across various existing copilots, from OpenAI’s models iteratively correcting their own generated code during sandbox execution to multi-step data-gathering workflows for complex tasks (Zheng et al., 2025).
In DECO, all iterative interactions are managed through the frontend, ensuring full transparency and user control. Past interactions are retained in history, allowing users to provide additional inputs when necessary. Unlike platforms that orchestrate multiple agents behind a single API call (see Figure 4(a)), DECO does not require custom handling mechanisms for multi-turn interactions but only execute one agent in one invocation of the backend (see Figure 4(b)), which offers better control compared to existing free-chat approaches, such as AutoGen (Wu et al., 2023).
(a) A single API call.
(b) Multi-turn interactions.
Figure 4. Frontend-backend interactions
Simplicity and Maintainability To facilitate ease of development, debugging, and extensibility, DECO adopts a modular and interpretable workflow based on a stepwise “tool-use scheme”. Execution follows a structured model where each agent sequentially invokes a subset of skills, actively interacting with the frontend. This modular execution allows engineers to inspect intermediate states, log agent outputs, and isolate failures efficiently. Since agents are loosely coupled, individual components can be updated, replaced, or extended without affecting the overall pipeline. Additionally, each execution round generates transparent logs, making it easier to trace and debug incorrect responses.
Low-Latency Execution DECO ensures real-time responsiveness by streaming intermediate outputs from agents as they execute. Unlike traditional batched execution, where users must wait for all skills to complete, our framework delivers partial responses incrementally (see Figure 4(b)). This approach enhances usability by allowing users to refine their input based on intermediate results, ensuring a dynamic and interactive experience. Furthermore, since each step displays output progressively, users can cancel an interaction at any time without losing already provided information.
Scalability and Efficient Planning Flat planners, which evaluate all skills simultaneously, become inefficient as the number of available skills grows. As the skill set expands, a flat planner must consider an exponentially increasing number of possible combinations, leading to computational inefficiencies and suboptimal execution sequences. Prior research, such as fine-tuning-based approaches (Patil et al., 2023), has attempted to mitigate this by training LLMs to generate skill execution plans. However, fine-tuning requires extensive labeled data, making it both costly and impractical in dynamic environments where new skills are frequently introduced.
To overcome these limitations, DECO employs a hierarchical agentic planning approach, dynamically selecting and sequencing agents based on the user query. Instead of treating skills as isolated decision points, the system:
  • First determines the relevant agent (e.g., troubleshooting, database analysis, code retrieval).
  • Then, based on the selected agent, determines the appropriate skills to invoke.
This hierarchical approach reduces complexity, improves efficiency, and maintains flexibility for integrating new skills, ensuring DECO remains adaptable as capabilities evolve.
5.2. Orchestration
Figure 5. Backend orchestration
Figure 5 illustrates the backend orchestration of the copilot, addressing the above consideration. In each round of backend workflow invocation, the Manager module performs the following functions:
  • Determines whether the previous invocation sufficiently answers the user’s question (Termination Decision Maker).
  • Dynamically updates the meta-plan, which specifies the sequence of agents to execute (Planner) and identifies the current agent to be executed in this round.
The Executor runs the current agent and generates the output. This output is sent to the frontend in streaming mode, automatically triggering the next workflow invocation (without user input) to execute the next agent in the meta-plan. This process continues until the Termination Decision Maker returns the termination flag. From the user interface, users can sequentially view outputs from each agent without waiting for the entire meta-plan execution to complete.
5.3. Agents and Skills
Agent
To enhance DECO’s modularity, we introduce the concept of “agents”, enabling onboarding teams to seamlessly integrate with various skills or even existing copilots through APIs. Each agent is defined by (1) a detailed description of applicable scenarios; (2) a set of skills the agent can invoke; (3) (optional) an LLM prompt that the agent uses at the end of execution to generate a response, combining all skill outputs as input context; and (4) (optional) a hint for the next agent to be invoked. Based on the setting of (3), the agent can either invoke an additional LLM call or directly return the skill output as the agent output for this invocation round.
For example, the Kusto Agent includes the skills kql.generator and kql.executor; the Root Cause Analysis Agent uses the RCA.copilot skill; the Code Agent utilizes the get_code skill; and the QA Agent employs the skills get_tsg, get_icm, etc.
Skill
Each skill extends from a common abstract class and is defined by (1) a detailed description of its functionalities; (2) a specification of its input arguments, such as a rephrased user question, which need to be extracted from the skill input; (3) the execution code (in Python), which may involve API calls, LLM interactions, or any other execution processes; and (4) (optional) skill dependencies, which impacts the execution order, ensuring that required outputs from preceding skills are available as inputs.
For example, the get_tsg skill requires a “rephrased user intent” as an argument and retrieves relevant documentation via the documentation search API. Its output consists of a list of TSG chunks, which can later be leveraged by the agent’s final LLM call to generate a response. The kql.executor skill can execute Kusto queries on behalf of users.
5.4. Manager
The Termination Decision Maker and the Planner execute in parallel during each workflow invocation round.
The Termination Decision Maker collects the chat history (i.e., the user’s question and responses from previously executed agents) and makes a termination decision using an LLM call. The prompt for this call includes detailed termination conditions and examples.
The Planner loads all available agents and compiles their names, descriptions, and hints for next agent into an agent selection prompt. An LLM generates the sequence of agents to be invoked in the meta-plan. In each round of workflow invocation, the Planner updates the meta-plan and selects the next agent for execution. In our current implementation, most user queries are resolved with fewer than three agents, and the Planner’s decisions are highly accurate.
5.5. Executor
Once the Planner selects an agent for execution, the Executor performs the following steps to “execute” the corresponding agent: Step 1: Selects the skills to execute from the agent’s available skill set and extract arguments; Step 2: Generates a skill execution plan based on defined skill dependencies (if any); Step 3: Executes the skill execution plan and collects outputs from all invoked skills; and Step 4 (Optional): Constructs the prompt based on the agent’s definition and makes a final LLM call to generate the agent’s response.
In Step 1, we compile descriptions of all available skills for the selected agent, along with their required input arguments, into the skill selector’s prompt. Few-shot examples are included to help guide the LLM in selecting the appropriate skills based on various user inputs. The output of this LLM call includes both the list of skills to invoke and the corresponding input arguments for each skill. For the skill orchestration LLM call, we use the Semantic Kernel (Microsoft, 2024g) sequential planner, integrated with Azure OpenAI function calling (Microsoft, 2024e) to restrict the output format. In Step 1, the LLM determines only the set of skills to execute. Step 2 then applies a deterministic optimization mechanism to decide the execution order and parallelism, accounting for dependency constraints. If no constraints exist, all skills are executed in parallel. For certain agents, some skills are capable of directly generating user responses and terminate this round of invocation.
Figure 6. NL2SearchQuery
6. Retrieval Algorithms
In this work we proposed lightweight retrieval strategies and re-ranking algorithms for skills such as get_icm, get_tsg, and get_code to enhance retrieval precision in production, supported by matured commercialized search services, such as Azure AI search (Microsoft, 2024b).
6.1. NL2SearchQuery
In the preprocessing pipeline, the processed data often includes several important search fields. For example, “resolve date” for IcMs, “title” and “content” for TSGs, and “reference” or “description” for code. Using an LLM, we can infer optimal matching criteria or additional filters based on user input to improve the retrieval results. For instance, queries such as “all the code related to the class SOSScheduler” or “show me recent incidents similar to Incident XXX” can be formulated as different search queries.
Since all fields are indexed, similar to a SQL query, we can search for relevant information by matching against different fields and applying different filters/predicates to pinpoint the most relevant documents. This insight informs the design of our NL2SearchQuery algorithm, as illustrated in Figure 6. Given a user query, we leverage an LLM call to extract key arguments for retrieval and construct a complete search query which consists of selected search fields, time ranges, or any other filtering criteria predefined for each retrieval skill. Multiple search queries can be executed in parallel, and the combined result set can be re-ranked using different ranking strategies to produce the final output. The following sections provide a detailed discussion of the retrieval strategies for IcM, TSG, and code respectively.
6.2. IcM Retrieval
@@
The get_icm skill, essential for generating curated incident reports, is frequently used to respond to user inquiries about how-to guides and to search for similar incidents, thereby offering additional mitigation strategies. We’ve noticed that user queries invoking this skill typically highlight specific incident details, such as failure mechanisms or server names, which align with the mitigation steps or extractable properties of incidents (discussed in Section 4). For this task, the following arguments are extracted to form the search query:
@@
  • Search Method: Options include hybrid search, semantic search, simple search, or vector search (Microsoft, 2024f).
@@
Re-ranking. The initial search query typically returns a large set of incidents, for example, 20. During the re-ranking step, we calculate scores for each incident to determine the top $K$ results, such as the top 4, for the final retrieval output.
The re-ranking score for a document $d$ is calculated as follows:
@@
where IS represents the information score, which evaluates the quality of the IcM summary based on data quality metrics like token length, pre-computed during the preprocessing pipeline (see Section 4). TS, the time score, assesses the relevance of incidents based on their age, with older incidents presumed less relevant and assigned lower values. SS, the source score, verifies if the retrieved data aligns with the current context, assigning a value of 1 for matches (e.g., matching team or monitor ID in property fields) and 0 otherwise. These scores are normalized and combined to re-rank the retrieved incidents, focusing on factors like information richness, timeliness, and source relevance. This process is detailed in Algorithm 1. The accuracy of our ICM retrieval algorithm significantly outperforms the default similarity search in the ICM Portal (see Section 8.1).
1:  Input: User question $q$, IcM retrieval database $D$, number of top results $K$, number of incidents to retrieve per query $N$, re-ranking score function $P(d)$
@@
7:     Compute the re-ranking score $P(d)$ for all $d\in D_{i}$
@@
10:  Rank all documents in $D^{*}$ based on the re-ranking score
@@
6.3. TSG Retrieval
The get_tsg function is a critical component of our backend flow. Traditional retrieval-augmented generation (RAG) approaches typically use similarity-based searches, which compare user query embeddings with document chunk embeddings or employ term-frequency techniques (Robertson et al., 2009). However, empirical analyses show that these methods often underperform with highly technical documents, as embeddings do not capture the subtle distinctions within technical terms effectively. Although recent advancements in supervised retrieval models, such as RAFT (Zhang et al., 2024b), AR2 (Zhang et al., 2022), and ARG (Sachan et al., 2023), have improved retrieval accuracy, they necessitate the development and fine-tuning of deep neural networks (DNNs). These approaches are impractical for our self-hosting framework in DECO, as they require specialized machine learning expertise and additional infrastructure for training and hosting ML models, which may not be readily available to different teams.
To address these challenges, we developed a hybrid search algorithm, leveraging an advanced preprocessing pipeline detailed in Section 4, eliminating the need for additional ML training while enhancing both retrieval accuracy and computational efficiency. Building on the NL2SearchQuery pipeline, we introduce multiple search queries that run in parallel, we extract the following arguments for the search query:
  • Rephrased Query: User queries are rephrased by the planner to generate search arguments. Searches are conducted using both the original user query and the generated search arguments. For search query involving incidents, we generate additional “search hint” based on the incident summary, similar to HyDE (Zhao et al., 2023; Gao et al., 2022).
  • Search Field: Both “title” and “contents”.
  • Search Method: Multi-Field Hybrid Search based on reciprocal rank fusion (RRF) (Cormack et al., 2009; Microsoft Corporation, 2024) is used to combine the similarity score from all the search fields.
We also introduce several additional techniques to enhance retrieval results. A new Document Filtering step refines search results by analyzing the similarity scores of top-ranked documents during the re-ranking stage. If the highest-ranked document(s) show a significant margin over the following results, only these top documents are retained. This approach leverages the observation that a small set of highly relevant TSGs is often sufficient to resolve issues, reducing the inclusion of overly detailed or irrelevant documents and saving tokens for the final LLM calls (see Section 8.1).
Additionally, during preprocessing, teams provide concise Repository Descriptions for each document source. These descriptions are included with the retrieved documents to enhance the LLM’s contextual understanding, improving the quality of the final LLM call within the agent.
Future enhancements include incorporating advanced retrieval algorithms like HyDE and Reverse HyDE (Gao et al., 2022; Zhao et al., 2023), generating enriched matching fields, and exploring the integration of graph structures into the index (Anonymous, 2024a) and leveraging feedback from historical conversations to dynamically refine search rankings (Anonymous, 2024b).
6.4. Code Retrieval
A similar retrieval mechanism is also employed in the get_code skill for code search based on LA-RAG using the NL2SearchQuery algorithm. With few-shot examples, the arguments learned based on the NL2SearchQuery module include: